OpenStudy (loser66):

Solve y"+y=cos^2(x) please, help

OpenStudy (loser66):

just x_p

OpenStudy (loser66):

by variation parameter

OpenStudy (reemii):

Does it help if you use \(\cos^2(x) = \frac{\cos(2x) + 1}{2}\)?

OpenStudy (loser66):

\[x_p=-\int sinx cos^2x dx *cosx +\int cos^3x dx sinx\]

OpenStudy (loser66):

for the first part, I let u = cos x, du = -sinx dx that gives me \(\dfrac{cos^3x}{3}\) .Am I right so far?

OpenStudy (loser66):

For the second part, the integrand becomes \(cos x (1-sin^2 x) = cos x -cosx sin^2x\) Hence \[\int( cosx -cosxsin^2x)dx *cos x\]

OpenStudy (loser66):

sorry, *sinx at the end

OpenStudy (loser66):

Hence it is \[sin^2 x -\int cosx sin^2x dx *sinx\]

OpenStudy (loser66):

Let u = sin x, du = cos x dx I got \(\dfrac{sin^4x}{3}\) All in

OpenStudy (loser66):

\[\dfrac{cos^4x}{3}-\dfrac{sin^4x}{3}+sin^2x\]

OpenStudy (loser66):

now, manipulate the first 2 terms. \((1/3) (cos^2 x)^2 -(sin^2 x)^2=(1/3) (cos^2 x-sin^2x)(cos^2x +sin^2x) \\=(1/3)( cos^2x -sin^2x)= (1/3)cos (2x)\)

OpenStudy (loser66):

so at the end, I got \((1/3)cos(2x) +sin^2x\) but wolfram gives me https://www.wolframalpha.com/input/?i=y%22%2By%3Dcos%5E2(x) why?

OpenStudy (reemii):

You are correct I guess. It's because of \(\cos(2a) = 2\cos^2(a) - 1 = \cos^2a-\sin^2a =1 - 2\sin^2(a)\). Therefore, \(\sin^2a = \frac{1-\cos(2a)}{2}\) and \(\frac13 \cos(2x) + \sin^2(x) = \frac13 \cos^2(x) + \frac12 - \frac12 \cos(2x) = \frac12 - \frac16 \cos(2x)\).

OpenStudy (loser66):

oh, so, just manipulate further and I will get what wolfram did, right? Thanks so much.

OpenStudy (reemii):

Yes, that happens often with the trigonometric functions. :-) Sometimes the "error" is hidden in the integrating function so that \(\cos^2(x)\) and \(2\cos^2(x)\) are "equal" (up to a constant).

OpenStudy (reemii):

integrating constant *

OpenStudy (reemii):

typo.. \(\cos(2x)\) and \(2\cos^2(x)\)

OpenStudy (loser66):

Got you. Thanks a lot.